A proton experiences an eastward force of 30 ×  N while traveling in a magnetic field of 5 ×  T north. 

  1. Find the magnitude of the proton’s velocity
  2. State the direction of the proton’s velocity 
  3. Draw the proton and indicate the directions of all three vectors using the appropriate notation as shown in the table above
1
Expert's answer
2021-05-03T10:29:26-0400

"F=qvB\sin \alpha=qvB,"

"v=\frac{F}{qB},"

"v=\frac{30}{1.6\cdot 10^{-19}\cdot 5}=37.5\cdot 10^{18}~\frac ms."